Malibu Corp. bond has a coupon rate of 8.25% (semiannual coupon) has a current yield of 5.25%. What is its yield to maturity? The bond has 4 years to maturity.
Answer #1

For computing the yield to maturity first we have to determine the present value by using the following formula

We assume the future value be $1,000

As we know that

Current yield = Annual coupon / Current price of the bond

5.25% = ($1,000 * 8.25%) / current price of the bond

5.25% = $82.50 / current price of the bond

So, the current price of the bond is $1,571.43

Now the yield to maturity could be found by using the RATE formula i.e to be shown in the spreadsheet below

Given that,

Present value = $1,571.43

Assuming figure - Future value or Face value = $1,000

PMT = 1,000 × 8.25% ÷ 2 = $41.25

NPER = 4 years × 2 = 8 years

The formula is shown below:

= -Rate(NPER,PMT,-PV,FV,type)

The present value come in negative

So, after applying the above formula

B5 fc =-RATE(B1,B2,-B3,B4) B С D E 8 e: А 1 NPER el 2 PMT 3 PV 4 FV 5 RATE n 6 41.25 1571.43 1000 2.2990%

The yield to maturity is

= 2.2990% * 2

= 4.60%